OpenStudy (anonymous):

A 42kg seal slides down a ramp into a pool. The top of the ramp is 1.75m higher than the surface of the water, and the ramp is inclined 35 degrees above the horizontal axis. If the seal reaches t

OpenStudy (anonymous):

the text is incomplete, but from what you typed i guess this is probably a problem of conservation of energy, in that case: \[mgh = \frac{ 1 }{ 2 }*m*v^2 => v=\sqrt{2gh} \\ d=\frac{ h }{ \sin \theta }\] v=velocity when the seal touches water d=distance before it its the water h=1.75 theta=35*pi/180 with some kinematics you can derive time taken as well but again the text was incomplete, so i probably answered in a wrong way!